8) The professor of a mechanical engineering course wishes to give each of 40 students a holiday gift. The amount needed to invest weekly for eight weeks at 26% nominal interest rate, compounded weekly,SO that each student will receive a $40 Jamba Juice gift card is nearest to a) $148 b) $164 c) $180 d) $198
Answer #1

Answer

  • Points to remember.
  • Amount (one of the option) is to be invested weekly.
  • Total Students = 40
  • Total amount needed for gift card after 8 weeks = 40 students x $ 40 card = $ 1,600
  • 26% is annual interest for 365 days.
    Per day Interest rate = 26/365 = 0.07%
    Weekly Interest rate = 0.07 x 7 days = 0.49%
  • CORRECT ANSWER = Option ‘D’ $ 198 is to be invested weekly.
  • Verification of answer:

Week #

Beginning balance of Investment

Amount invested

Total Investment

Interest at 0.49%

Ending balance of Investment

[A]

[B]

[C = A + B]

[D = C x 0.49%]

[E = C + D]

1

$                                               -  

$                      198

$                     198

$                   0.97

$                     199

2

$                                            199

$                      198

$                     397

$                   1.95

$                     399

3

$                                            399

$                      198

$                     597

$                   2.92

$                     600

4

$                                            600

$                      198

$                     798

$                   3.91

$                     802

5

$                                            802

$                      198

$                 1,000

$                   4.90

$                 1,005

6

$                                        1,005

$                      198

$                 1,203

$                   5.89

$                 1,209

7

$                                        1,209

$                      198

$                 1,407

$                   6.89

$                 1,413

8

$                                        1,413

$                      198

$                 1,611

$                   7.90

$                 1,619

--Investing $ 198 weekly will give $ 1619 aprrox at the end of 8 week period, which is near around the money needed to provide for 40 students gift card at $ 40 each [$ 1600]